Circular No. 8084

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003bh AND 2003bi Further to IAUC 8082, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of two supernovae on Palomar NEAT images.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003bh Feb. 21.32 10^{h}00^{m}06.^{s}27 +28^{o}16′52′′.0 19.9 7′′.5 W, 2′′.0 N
2003bi Feb. 22.43 10 44 05.97 +12 31 34.7 19.6 2′′.3 E, 3′′.1 N

Additional magnitudes from NEAT images: SN 2003bh, Jan. 10.52 UT, [20.7 (S/N = 3). SN 2003bi, Jan. 28.52, [20.4 (S/N = 3); Feb. 23.49, 19.6.

A. V. Filippenko and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Feb. 27 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2003bh is of type Ia, ∼ 1 month past maximum brightness; its host galaxy has a redshift of 0.089. A spectrum from Feb. 28 shows that SN 2003bi is also of type Ia, close to maximum; its host galaxy has a redshift of 0.093.

NO SUPERNOVA 2003aw Chornock and Filippenko add that inspection of CCD spectra, ob-

tained as above on Feb. 28 UT, shows that the object 2003aw (IAUC 8077) is not a supernova, but rather is probably a hydrogen-deficient dwarf nova similar to that proposed for 1998di = KL Dra on IAUC 6983. The spectrum consists of a nearly featureless, quite blue continuum. Very-low-contrast He I lines are present near zero redshift in emission, and there is Ca II H and K absorption at essentially zero redshift; they are too broad to be interstellar, and no interstellar Na I D absorption is visible.

SUPERNOVAE 2003aa AND 2003bg Filippenko and Chornock also report that inspection of CCD spectra,

obtained as above on Feb. 27 or 28 UT, shows that SN 2003aa (IAUC 8063) appears to be a type-Ia supernova similar to the subluminous SN 1991bg (e.g., Filippenko et al. 1992, A.J. 104, 1543), a few weeks past maximum brightness, rather than a type-Ic supernova as suggested in IAUC 8064. However, the expansion velocity of the ejecta is very low, resembling that of SN 2002es (IAUC 7965). SN 2003bg (IAUC 8082) is very similar to the peculiar type-Ic supernovae (called ‘hypernovae’ by some workers) 1997ef, 1998bw, 2002ap, and 2002bl, perhaps 2 weeks after the explosion (Iwamoto et al. 2000, Ap.J. 534, 660, Fig. 1).

2003 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green